OpenStudy (danysanchy97):

Which one is a proper fraction? A. 11/10 B. 6/6 C. 4/17 D. 15/2

OpenStudy (imsleeziboii):

4/17

OpenStudy (solomonzelman):

Proper fraction is: \(\color{#000000}{ \displaystyle \frac{a}{b} }\) where \(\color{#000000}{ \displaystyle b>a }\)

OpenStudy (imsleeziboii):

A is improper B is equal to 1 C is proper D is improper

OpenStudy (solomonzelman):

So, for example: \(\color{#000000}{ \displaystyle \frac{1}{2} ,~~ \frac{12}{127} ,~~ \frac{5}{67} ,~~ \frac{48}{49} ,~~\frac{5}{6} }\)

OpenStudy (solomonzelman):

those would be proper fractions

OpenStudy (imsleeziboii):

a proper fraction means the numerator (top #) has to be smaller than the demoninator(bottom number)

OpenStudy (imsleeziboii):

medal and fan if you need more help :)

OpenStudy (shaswatrajput):

A proper fraction is in which numerator is less than denominator or we can say that it is less than 1